Piyush spends his entire income between food grains and books (both of which are normal


goods) Pandemic causes a large increase in the prices of food grains in India. i. Show the effect of Pandemic on Piyush's budget constraint.


ii.


Show the effect of Pandemic on Piyush's optimal consumption


substitution effect outweighs the income effect for books.


bundle assuming that


iii. Show the effect of Pandemic on Piyush's optimal consumption bundle assuming


income effect outweighs the substitution effect for books.


that


1
Expert's answer
2021-03-26T08:06:57-0400

(I)Books and grains are normal goods meaning an increase in price of grains wi lead to increased quantity demanded for books than grains.

Piyush's budget constraint before pandemic was originally

"(PF*QF)+(PB*QB)=Y"

PF=price of grains

QF=quantity of food grains consumed

PB=price of books

QB=quantity of books consumed.

After increase in price of food grains, the price of food grains increases from PF to PF1 and the quantity decreases from QF to QF1. The new budget constraint will be

"(PF1*QF1)+(PB*QB)=Y"

(ii) substitution effect will increase the quantity demanded for books from QB to QB1)

"(PF1*QF1)+(PB*QB1)=Y"

(iii) Income effect increases the income from Y to Y1

"(PF1*QF1)+(PB*QB1)=Y1"